# Summarization Metrics for Spanish and Basque: Do Automatic Scores and LLM-Judges Correlate with Humans?

### v1

@@ 3.3.2 Annotation Procedure
We performed three rounds of expert annotation, where the first two served to refine the annotation guidelines. Given the generally poor agreement of crowdsourced annotations for summarization (?, ?), the annotators were recruited among NLP linguists and engineers who had extensive experience with annotation projects and knowledge of summarization.
@@
After each round of annotation, we calculated interannotator agreement using ordinal Krippendorff’s $\alpha$ (?) to compare three-way agreement for each of the criteria, and additionally calculated quadratic Cohen’s $\kappa$ (?) for pairwise agreement. Results are discussed in the next section.
@@ 4.1 Interannotator Agreement
The Krippendorff’s $\alpha$ correlations for each round and language are shown in Table 2. Initial annotations (R#=1) yielded moderate agreement levels for Basque across criteria and lower agreement for Spanish. By the final round (R#=3), Basque annotations maintained acceptable agreement levels (0.44–0.72), with Spanish remaining at lower values (0.19–0.39). Consistent with prior work (?, ?), consistency proved to be the most challenging for annotators to align on—0.44 and 0.19 for Basque and Spanish, respectively, in the final round. This challenge can be attributed to the inherent complexity of evaluating factual alignment, where annotator’s interpretations of source article details sometimes diverge.
Additional analyses presented in Appendix C.1 provide further insights into agreement patterns. While pairwise quadratic Cohen’s $\kappa$ values follow similar patterns to the Krippendorff’s $\alpha$ scores, raw agreement percentages show moderate to high concordance in several categories for Spanish, notably fluency (83–90%) and consistency (65–67%). This apparent contradiction is explained by the annotation distributions: in contrast to Basque summaries, which exhibit a broader distribution across the quality spectrum, Spanish summaries tend to concentrate in the higher range of the quality scale, leading to higher expected agreement and consequently lower chance-corrected scores.
Note that our agreement scores, including those for Spanish annotations, are similar to or exceed those reported in previous large-scale summary evaluation studies of reference (?, ?, ?).
@@
Table 2: Ordinal Krippendorff’s $\alpha$ agreement between annotators across triple-annotated rounds (R#). Rounds 1 and 2 were conducted on the same set of summaries (Doc), before and after guideline discussion respectively, while round 3 involved a new set of summaries.
4.2 Quantitative Analysis
We analyze the diversity of summarization strategies through two key metrics. Compression ratio indicates summary conciseness with respect to the source article (?), while novel 2-grams measure the percentage of new word pairs in the summary (?).
@@ 4.2 Quantitative Analysis
Table 3: Statistics of the BASSE corpus, including the source articles and their summaries. We report the number of documents (Doc)—and, of those, how many are not in the target language (¬ eu or ¬ es)—, total tokens (Tok), vocabulary size (Voc), and sentences per document (s/d). Splitting and tokenization were done with stanza (?) and language identification with lingua-py. [^12] https://github.com/pemistahl/lingua-py
@@ 4.3 Qualitative Analysis
The manual evaluation of the summaries (detailed results available in Table 4, Appendix C.2) shows significant quality differences across summarization approaches. While human-written summaries achieve the highest average scores (4.75 for Basque, 4.68 for Spanish), that is not the case for the subheads, which fail at covering key information (criterion 5W1H).
@@ 5 Evaluation
In this section, we calculate the correlation between the human annotations from Section 4 and a series of evaluation metrics and LLM-as-a-Judge commonly used for evaluating NLG. For rounds with more than one annotation per summary, we take the average of the three values for each criterion. We calculate both Spearman’s ($\rho$) and Kendall’s ($\tau$) rank correlation coefficients at system-level following ? (?) and ? (?), but report only Spearman in the subsequent section, as the results do not vary depending on the coefficient.
@@ 5.1 Evaluation Metrics
We select a subset of metrics used in SummEval (?) which require minimal changes to work for languages other than English. Due to their dependence on specialized English training data or English-only models, several of the original metrics are excluded, namely, BLANC (?), S3 (?), SummaQA (?), MoverScore (?), Sentence Mover’s Similarity (?), and SUPERT (?). For several other metrics (i.a., ROUGE-we, BertScore, METEOR, CIDEr, and Data Statistics, explained below), we implement the necessary changes to allow for evaluation in Spanish or Basque. Before passing the candidate or reference summaries to each metric, we perform metric specific preprocessing of the texts (see details in the description of each metric).
@@
  • ROUGE (?) is a series of metrics that calculate token subsequence overlap between a candidate summary and a set of reference summaries. We calculate ROUGE-{1–4}, ROUGE-L, and ROUGE-su* variants.
@@
  • ROUGE-we (?) extends the original ROUGE metric by incorporating soft matching based on cosine similarity of word embeddings. We use FastText embeddings for both Basque and Spanish. [^13] https://fasttext.cc/docs/en/crawl-vectors
@@
  • BertScore (?) calculates the similarity of two texts as the average token-level cosine similarities between greedily aligned tokens in each text. For our experiments, we replace BERT with multilingual BERT base uncased [^14] https://huggingface.co/google-bert/bert-base-multilingual-uncased and report precision (p), recall (r), and F1 (f) scores.
@@
  • BLEU (?) is a corpus-level metric originally designed for machine translation evaluation. It calculates average n-gram precision, coupled with a brevity penalty. We specifically use the SacreBLEU implementation (?).
@@
  • CHRF (?) calculates character-level n-gram overlaps between the candidate summary and the reference summary.
@@
  • METEOR (?) computes an alignment between candidate and reference sentences, mapping unigrams in the generated summary to 0 or 1 unigrams in the reference, based on stems, synonyms, and paraphrase matches. Precision and recall are computed and reported as a harmonic mean. We compute Multilingual METEOR (mMETEOR) by setting the language to other and only lower-casing the input.
@@
  • CIDEr (?) is a corpus-level metric originally designed for image description evaluation. It calculates the average cosine similarity between TD-IDF weighted n-gram representations of the candidate summary and a set of reference summaries, taking the average score of 1–4 grams. As this metric applies stemming, we use the Basque and Spanish versions of the Snowball Stemmer (?).
@@
  • Data Statistics (?) calculates a series of metrics originally designed to describe how extractive a dataset is. This includes i) Coverage, which measures how derivative a summary is; ii) Density, which measures how well a summary can be described as a series of extractions from the original text or iii) Compression, which is the ratio of the length of the summary divided by the length of the original text. SummEval further introduces n-gram related novelty—novel {1–3} gram—and redundancy scores—repeated {1–3} gram. Finally, we also include the length of the generated summaries.
@@ 5.2 LLM-as-a-Judge
  • Prometheus 2 (?) is an open-source LLM trained to evaluate the output of other models following an annotation rubric. Although Prometheus 2 is primarily trained for English, its backbone model Mistral-7B-Instruct-v0.2151515https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 is a multilingual model able to process Spanish, but not Basque. We experiment with both the 7B and 8x7B parameter models.
  • Selene Mini (?) is similarly an open-source LLM [^16] https://huggingface.co/AtlaAI/Selene-1-Mini-Llama-3.1-8B that uses a mix of reinforcement learning and supervised fine-tuning to train Llama3.1-8B-Instruct (?) to act as a general-purpose evaluator. For English, this model shows better performance than previous open-source models and is comparable to small closed-source models.
  • Qwen2.5 (?) is a set of open instruction-tuned LLMs that have shown promise as judge models in English (?). It officially supports Spanish, and authors report transfer to low-resource languages through translated data. We specifically use the 7B and 72B instruction-tuned models.
  • GPT-4o is a general-purpose commercial LLM that has shown strong performance as a judge model in English (?). It has also shown robust performance in other Spanish and Basque tasks (?, ?, ?). We experiment with GPT-4o mini and GPT-4o. [^17] The specific models used were gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06.
We use the codebase from Prometheus-Eval [^18] https://github.com/prometheus-eval/prometheus-eval and format the prompt to include the original instruction, the annotation rubric for the criteria, a human-generated reference answer, and the summary to be evaluated. We evaluate only one criterion at a time. Although the original documents and summaries are in either Basque or Spanish, we use English for the instructions and annotation rubric, as preliminary results indicated that several models struggle to follow the directions otherwise. The full prompt template used for all judge models is shown in Appendix D.
@@ 6 Results
In this section, we answer the research questions by analyzing Spearman’s $\rho$ correlations between model rankings derived from automatic evaluations and those from human judgments. Selected results are shown in Figure 2, while complete results of both Spearman’s $\rho$ and Kendall’s $\tau$ rank correlation coefficients can be consulted in Appendix C.3.
@@
Relevance has relatively high absolute mean correlations (0.501/0.358). Statistics-based metrics have generally high correlations in both languages—e.g., length (-0.887/-0.621), novel unigram (-0.780/-0.686), novel bigram (-0.756/-0.484)—, while BertScore-p (0.737/0.801) and CIDEr (0.677/0.654) also capture this criterion well across both languages. Most ROUGE variants, however, perform poorly, as do all LLM-as-a-Judge models.
@@
The open-source LLM-as-a-Judge models, on the other hand, do not generally outperform the automatic metrics, with the exception of Selene Mini with $\rho$ of 0.812 for 5W1H in Basque. Both Prometheus 2 8x7B and Qwen2.5-72B-Instruct perform as poorly as their smaller versions, suggesting that the low correlation achieved is not due to model size. Furthermore, both Prometheus 2 judges show better—though still very weak—correlations in Basque than in Spanish, which is particularly notable given that the backbone model of Prometheus—Mistral—officially supports Spanish but not Basque.
@@
Overall, for the models tested, coherence is the most problematic criterion currently in both languages, while fluency and relevance may be challenging specifically in Basque. Detailed results are available in Table 4 (Appendix C.2).
@@ 7.1 Methodological Considerations
Although our corpus collection and annotation methodology are the same for both Basque and Spanish, the outcome differs between languages. In Basque, the summaries we collected are generally of poorer quality, but the IAA is high, while for Spanish the quality of all summaries is quite good, but the IAA is lower. This seems to be an artifact of the higher quality of the summaries in Spanish. Annotators rarely used the lower end of the scale (1–3) for several criteria, and the expected agreement is therefore higher, affecting IAA scores, despite raw agreement often being over 80% (see Appendix C.1).
@@
Finally, note that most of the news articles in our dataset have a single human-generated summary, which limits our ability to exploit the variability of human summaries. However, given a set annotation budget and the goal of creating a large and varied set of summaries, previous research has shown it is preferable to increase the number of instances, even if they are single annotated (?).
@@ 7.2 On Contributions and Findings
Finally, we observed that—despite LLMs being capable of producing mostly faithful summaries—consistency is currently poorly captured by LLM-as-a-Judge models, as well as by all automatic metrics, contrary to previous findings in English (?). This is likely due to the choice of summarization models, as previous meta-evaluations have included summaries from extractive summarization models and simpler abstractive models. Current abstractive summarization models often have subtle inconsistencies (?, ?), potentially making them more difficult to capture.

### v2

@@ 3.3.2 Annotation Procedure
We performed three rounds of expert annotation, where the first two served to refine the annotation guidelines. Given the generally poor agreement of crowdsourced annotations for summarization Gillick and Liu (2010); Fabbri et al. (2021), the annotators were recruited among NLP linguists and engineers who had extensive experience with annotation projects and knowledge of summarization.
@@
After each round of annotation, we calculated interannotator agreement using ordinal Krippendorff’s $\alpha$ Krippendorff (2013) to compare three-way agreement for each of the criteria, and additionally calculated quadratic Cohen’s $\kappa$ Cohen (1960) for pairwise agreement. Results are discussed in the next section.
@@ 4.1 Interannotator Agreement
The Krippendorff’s $\alpha$ correlations for each round and language are shown in Table 2. Initial annotations (R#1) yielded moderate agreement levels for Basque across criteria and lower agreement for Spanish. By the final round (R#3), Basque annotations maintained acceptable agreement levels (0.44–0.72), with Spanish remaining at lower values (0.19–0.39). Consistent with prior work Falke et al. (2019); Kryscinski et al. (2020), consistency proved to be the most challenging for annotators to align on—0.44 and 0.19 for Basque and Spanish, respectively, in the final round. This challenge can be attributed to the inherent complexity of evaluating factual alignment, where annotator’s interpretations of source article details sometimes diverge.
While pairwise quadratic Cohen’s $\kappa$ values (Appendix C) follow similar patterns to the Krippendorff’s $\alpha$ scores, raw agreement percentages (Figures 5 and 5) show moderate to high concordance in several categories for Spanish, notably fluency (83–90%) and consistency (65–67%). This apparent contradiction is explained by the score distributions, depicted in Figures 5 and 5: in contrast to Basque summaries, which exhibit a broader distribution across the quality spectrum, Spanish summaries tend to concentrate in the higher range of the quality scale, leading to higher expected agreement and consequently lower chance-corrected scores.
Note that our agreement scores, including those for Spanish annotations, are similar to or exceed those reported in previous large-scale summary evaluation studies of reference Fabbri et al. (2021); Clark et al. (2023); Aharoni et al. (2023).
4.2 Quantitative Analysis
We analyze the diversity of summarization strategies through two key metrics. Compression ratio indicates summary conciseness with respect to the source article Segarra Soriano et al. (2022), while novel 2-grams measure the percentage of new word pairs in the summary Kryściński et al. (2018).
Table 2: Ordinal Krippendorff’s $\alpha$ across triple-annotated rounds (R#). Rounds 1 and 2 were conducted on the same set of summaries (Doc), before and after guideline discussion respectively, while round 3 involved a new set of summaries.
@@ 4.2 Quantitative Analysis
Figure 2: Agreement percentage in Basque R#3
Figure 3: Agreement percentage in Spanish R#3
Figure 4: Score distribution across annotators in Basque R#3
Figure 5: Score distribution across in Spanish R#3
Table 3: Statistics of the BASSE corpus, including the source articles and their summaries. We report the number of documents (Doc)—and, of those, how many are not in the target language (¬ eu or ¬ es)—, total tokens (Tok), vocabulary size (Voc), and sentences per document (s/d). Splitting and tokenization were done with stanza Qi et al. (2020) and language identification with lingua-py. [^14] https://github.com/pemistahl/lingua-py
@@ 4.3 Qualitative Analysis
{tblr}
[…]
Human 4.77 4.94 4.85 4.51 4.69 4.75 4.91 4.88 4.80 4.48 4.34 4.68
LLM 3.46 4.29 3.87 3.74 4.06 3.89 3.88 4.68 4.82 4.14 4.18 4.34
by model
  Claude 3.17 4.61 3.83 3.54 4.57 3.95 3.17 4.71 4.40 3.81 4.40 4.10
  GPT-4o 3.88 4.52 4.55 3.78 4.62 4.27 4.13 4.77 4.94 4.37 4.37 4.52
  Reka 3.72 4.20 4.05 4.02 4.14 4.03 4.21 4.49 4.87 4.26 4.17 4.40
  LLama 3.1 3.59 4.50 4.29 4.31 3.45 4.03 3.78 4.79 4.93 3.98 3.94 4.29
  Cmd R+ 2.95 3.65 2.64 3.07 3.53 3.17 4.10 4.66 4.94 4.29 4.02 4.40
by prompt
  Base 3.78 4.20 3.70 3.36 4.19 3.85 4.22 4.67 4.91 4.12 4.20 4.42
  Core 3.61 4.33 4.05 3.73 3.64 3.87 4.11 4.73 4.93 4.08 3.94 4.36
  5W1H 2.62 4.22 4.10 3.64 4.39 3.79 3.03 4.58 4.96 4.03 4.65 4.25
  tldr 3.82 4.42 3.64 4.24 4.01 4.03 4.14 4.75 4.47 4.35 3.94 4.33
[…]
The manual evaluation of the summaries (see Table 4; detailed results per model and prompt combination are available in Appendix D) shows significant quality differences across summarization approaches. While human-written summaries achieve the highest average scores (4.75 for Basque, 4.68 for Spanish), that is not the case for the subheads, which fail at covering key information (criterion 5W1H).
@@ 5 Evaluation
In this section, we calculate the correlation between the human annotations from Section 4 and a series of evaluation metrics and LLM-as-a-Judge commonly used for evaluating NLG. For rounds with more than one annotation per summary, we take the average of the three values for each criterion. We calculate both Spearman’s ($\rho$) and Kendall’s ($\tau$) rank correlation coefficients at system-level following Louis and Nenkova (2013) and Fabbri et al. (2021), but report only Spearman in the subsequent section, as the results do not vary depending on the coefficient.
@@ 5.1 Evaluation Metrics
We select a subset of metrics used in SummEval Fabbri et al. (2021) which require minimal changes to work for languages other than English. Due to their dependence on specialized English training data or English-only models, several of the original metrics are excluded, namely, BLANC Vasilyev, Dharnidharka, and Bohannon (2020), S3 Peyrard, Botschen, and Gurevych (2017), SummaQA Scialom et al. (2019), MoverScore Zhao et al. (2019), Sentence Mover’s Similarity Clark, Celikyilmaz, and Smith (2019), and SUPERT Gao, Zhao, and Eger (2020). For several other metrics (i.a., ROUGE-we, BertScore, METEOR, CIDEr, and Data Statistics, explained below), we implement the necessary changes to allow for evaluation in Spanish or Basque. Before passing the candidate or reference summaries to each metric, we perform metric specific preprocessing of the texts (see details in the description of each metric).
@@
  • ROUGE Lin (2004) is a series of metrics that calculate token subsequence overlap between a candidate summary and a set of reference summaries. We calculate ROUGE-{1–4}, ROUGE-L, and ROUGE-su* variants.
@@
  • ROUGE-we Ng and Abrecht (2015) extends the original ROUGE metric by incorporating soft matching based on cosine similarity of word embeddings. We use FastText embeddings for both Basque and Spanish. [^15] https://fasttext.cc/docs/en/crawl-vectors
@@
  • BertScore Zhang et al. (2020) calculates the similarity of two texts as the average token-level cosine similarities between greedily aligned tokens in each text. For our experiments, we replace BERT with multilingual BERT base uncased [^16] https://huggingface.co/google-bert/bert-base-multilingual-uncased and report precision (p), recall (r), and F1 (f) scores.
@@
  • BLEU Papineni et al. (2002) is a corpus-level metric originally designed for machine translation evaluation. It calculates average n-gram precision, coupled with a brevity penalty. We specifically use the SacreBLEU implementation Post (2018).
@@
  • CHRF Popović (2017) calculates character-level n-gram overlaps between the candidate summary and the reference summary.
@@
  • METEOR Banerjee and Lavie (2005) computes an alignment between candidate and reference sentences, mapping unigrams in the generated summary to 0 or 1 unigrams in the reference, based on stems, synonyms, and paraphrase matches. Precision and recall are computed and reported as a harmonic mean. We compute Multilingual METEOR (mMETEOR) by setting the language to other and only lower-casing the input.
@@
  • CIDEr Vedantam, Zitnick, and Parikh (2015) is a corpus-level metric originally designed for image description evaluation. It calculates the average cosine similarity between TD-IDF weighted n-gram representations of the candidate summary and a set of reference summaries, taking the average score of 1–4 grams. As this metric applies stemming, we use the Basque and Spanish versions of the Snowball Stemmer Porter (2001).
@@
  • Data Statistics Grusky, Naaman, and Artzi (2018) calculates a series of metrics originally designed to describe how extractive a dataset is. This includes i) Coverage, which measures how derivative a summary is; ii) Density, which measures how well a summary can be described as a series of extractions from the original text or iii) Compression, which is the ratio of the length of the summary divided by the length of the original text. SummEval further introduces n-gram related novelty—novel {1–3} gram—and redundancy scores—repeated {1–3} gram. Finally, we also include the length of the generated summaries.
@@ 5.2 LLM-as-a-Judge
  • Prometheus 2 Kim et al. (2024) is an open-source LLM trained to evaluate the output of other models following an annotation rubric. Although Prometheus 2 is primarily trained for English, its backbone model Mistral-7B-Instruct-v0.2171717https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2 is a multilingual model able to process Spanish, but not Basque. We experiment with both the 7B and 8x7B parameter models.
  • Selene Mini Alexandru et al. (2025) is similarly an open-source LLM [^18] https://huggingface.co/AtlaAI/Selene-1-Mini-Llama-3.1-8B that uses a mix of reinforcement learning and supervised fine-tuning to train Llama3.1-8B-Instruct Dubey et al. (2024) to act as a general-purpose evaluator. For English, this model shows better performance than previous open-source models and is comparable to small closed-source models.
  • Qwen2.5 Qwen et al. (2025) is a set of open instruction-tuned LLMs that have shown promise as judge models in English Gu et al. (2025). It officially supports Spanish, and authors report transfer to low-resource languages through translated data. We specifically use the 7B and 72B instruction-tuned models.
  • GPT-4o is a general-purpose commercial LLM that has shown strong performance as a judge model in English Huang et al. (2024). It has also shown robust performance in other Spanish and Basque tasks Achiam et al. (2023); Etxaniz et al. (2024a, b). We experiment with GPT-4o mini and GPT-4o. [^19] The specific models used were gpt-4o-mini-2024-07-18 and gpt-4o-2024-08-06.
We use the codebase from Prometheus-Eval [^20] https://github.com/prometheus-eval/prometheus-eval and format the prompt to include the original instruction, the annotation rubric for the criteria, a human-generated reference answer, and the summary to be evaluated. We evaluate only one criterion at a time. Although the original documents and summaries are in either Basque or Spanish, we use English for the instructions and annotation rubric, as preliminary results indicated that several models struggle to follow the directions otherwise. The full prompt template used for all judge models is shown in Figure 7.
###Task Description:
###The instruction to evaluate:
{instruction}
###Response to evaluate:
{response}
###Reference Answer (Score 5):
{reference answer}
###Score Rubrics:
{rubric}
###Feedback:
Figure 7: Prompt from Prometheus-Eval used for LLM-as-a-Judge evaluation. Note that instruction, response, reference answer, and rubric are placeholders that are filled with specific information for each example.
@@ 6 Results
In this section, we answer the research questions by analyzing Spearman’s $\rho$ correlations between model rankings derived from automatic evaluations and those from human judgments. Selected results are shown in Figure 8, while full results of both Spearman’s $\rho$ and Kendall’s $\tau$ rank correlation coefficients can be consulted in Appendix E.
@@
Relevance has relatively high absolute mean correlations (0.501/0.358). Statistics-based metrics have generally high correlations in both languages—e.g., length (-0.887/-0.621), novel unigram (-0.780/-0.686), novel bigram (-0.756/-0.484)—, while BertScore-p (0.737/0.801) and CIDEr (0.677/0.654) also capture this criterion well across both languages. Most ROUGE variants, however, perform poorly, as do all judge models.
@@
RQ2: Do LLM-as-a-judge models have the ability to replicate human judgments for summarization?
@@
The open-source LLM-as-a-Judge models, on the other hand, do not generally outperform the automatic metrics, with the exception of Selene Mini with $\rho$ of 0.812 for 5W1H in Basque. Both Prometheus 2 8x7B and Qwen2.5-72B-Instruct perform as poorly as their smaller versions, suggesting that the low correlation achieved is not due to model size. Furthermore, both Prometheus 2 judges show better—though still very weak—correlations in Basque than in Spanish, which is particularly notable given that the backbone model of Prometheus, Mistral, officially supports Spanish but not Basque.
@@
Overall, for the models tested, coherence is the most problematic criterion currently in both languages, while fluency and relevance may be challenging specifically in Basque. Detailed results are available in Appendix D.
@@ 7.1 Methodological Considerations
Although our corpus collection and annotation methodology are the same for both Basque and Spanish, the outcome differs between languages. In Basque, the summaries we collected are generally of poorer quality, but the IAA is high, while for Spanish the quality of all summaries is quite good, but the IAA is lower. This seems to be an artifact of the higher quality of the summaries in Spanish. Annotators rarely used the lower end of the scale (1–3) for several criteria, and the expected agreement is therefore higher, affecting IAA scores, despite raw agreement often being over 80%.
@@
Finally, note that most of the news articles in our dataset have a single human-generated summary, which limits our ability to exploit the variability of human summaries. However, given a set annotation budget and the goal of creating a large and varied set of summaries, previous research has shown it is preferable to increase the number of instances, even if they are single annotated Steen and Markert (2021).
@@ 7.2 On Contributions and Findings
Finally, we observed that—despite LLMs being capable of producing mostly faithful summaries—consistency is currently poorly captured by LLM-as-a-Judge models, as well as by all automatic metrics, contrary to previous findings in English Fabbri et al. (2021). This is likely due to the choice of summarization models, as previous meta-evaluations have included summaries from extractive summarization models and simpler abstractive models. Current abstractive summarization models often have subtle inconsistencies Kryscinski et al. (2020); Yang et al. (2024), potentially making them more difficult to capture.